OpenStudy (anonymous):

12.5% of what number is 130

Parth (parthkohli):

\[12.5\% \times x = 130 \\ \implies {12.5 \over 100} x = 130 \\ \implies 0.125x = 130 \]

OpenStudy (anonymous):

Let the number be x

OpenStudy (amistre64):

take a value: n 1 percent of n is just: n/100 12.5 of those 1 percents is: 12.5n/100 and that needs to equate to 130\[\frac{12.5n}{100}=130\]

OpenStudy (anonymous):

12.5% = 130 Divide both sides by 12.5 to get 1%: 1% = 130/12.5 = 10.4 Multiply both sides by 100 to get 100%: 100% = 1040